OpenStudy (anonymous):

integral of cos(3x-5)^2

OpenStudy (anonymous):

use double angles rules

OpenStudy (anonymous):

cos2x=2cosx-1 .. did you get it ?

OpenStudy (anonymous):

and what about the cos^2

OpenStudy (anonymous):

sry cos2x=2cosx^2-1

myininaya (myininaya):

Recall the double angle identity for cosine. \[\cos(2 \alpha )=\cos^2(\alpha)-\sin^2(\alpha)=\cos^2(\alpha)-(1-\cos^2(\alpha))\] \[\cos(2 \alpha)=\cos^2(\alpha)-1+\cos^2(\alpha)\] \[\cos(2 \alpha)=2 \cos^2(\alpha)-1\] Solve this equation for cos^2(alpha): \[\cos^2(\alpha)=\frac{1}{2}(1+\cos(2 \alpha))\]

OpenStudy (anonymous):

Thank you Amrmagdy and Myininaya :D

OpenStudy (anonymous):

urw